OpenStudy (anonymous):

how can we find determinant(5)=? explain its procedure

OpenStudy (anonymous):

what is the matrix?

OpenStudy (zarkon):

Det([5])=5

OpenStudy (anonymous):

what is the method of calculating it

OpenStudy (anonymous):

we have method of finding 2x2 and 3x3 matrices but how can we say that det([5])=5 tell me the logic

OpenStudy (zarkon):

we would still like to retain some of the properties of determinates Det(AB)=Det(A)Det(B) when A,B are square and Det(identity)=1 so it is natural to define Det([a])=a

OpenStudy (anonymous):

you mean there is no logic

OpenStudy (zarkon):

for any square (nxn) matrix A \[Det(c\cdot A)=c^nDet(A)\] \[Det([a])=Det(a\cdot[1])=a^1Det([1])=a\cdot 1=a\]

OpenStudy (anonymous):

oh yes this is a proper answer thanks a lot.